# Topological and nontopological degeneracies in generalized string-net models

### v3

@@ Abstract
Generalized string-net models have been proposed recently in order to enlarge the set of possible topological quantum phases emerging from the original string-net construction. In the present work, we do not consider vertex excitations and we restrict ourselves to plaquette excitations, or fluxons, that satisfy important identities. We explain how to compute the energy-level degeneracies of the generalized string-net Hamiltonian associated with an arbitrary unitary fusion category. In contrast to the degeneracy of the ground state, which is purely topological, the degeneracy of excited energy levels depends not only on the Drinfeld center of the category, but also on internal multiplicities obtained from the tube algebra defined from the category. For a noncommutative category, these internal multiplicities result in extra nontopological degeneracies. Our results are valid for any trivalent graph and any orientable surface. We illustrate our findings with nontrivial examples.
@@ I Introduction
The idea that two-dimensional quantum systems could contain quasiparticles with exchange statistics different from bosons and fermions emerged about 50 years ago [1, 2]. These point-like particles, called anyons, could, under mutual exchange, accumulate “any,” possibly fractional, phase (Abelian anyons), or even change into a new state orthogonal to the initial one (non-Abelian anyons). Yet, only recent experiments provided evidence for the presence of Abelian anyons in the fractional quantum Hall effect [3, 4], and investigated non-Abelian anyons in quantum processors made of superconducting qubits [5, 6] or trapped ions [7]. Anyons are a fundamental characteristic of topologically ordered phases: gapped phases that appear in strongly interacting or frustrated systems and that cannot be identified by local symmetries. In particular, anyons are connected to one of the defining properties of topological order: a ground-state degeneracy, which depends on the surface topology on which the system is defined, and which is robust against small local perturbations (a fact that motivated the idea of using topological order to implement fault-tolerant quantum computation  [8, 9]). The excited states of topologically ordered phases also show non-trivial degeneracies related to the fusion of anyons.
Some of the most studied models for topologically ordered phases are the string-net models introduced by Levin and Wen [10], and their generalizations [11, 12, 13, 14]. These models, while not being able to produce all possible topological orders, are believed to produce exactly those bosonic topological orders that have gappable edges (and are therefore achiral). The string-net models are exactly solvable. While the energy levels are easy to obtain, the corresponding spectrum degeneracies are non-trivial and are the main focus of the present paper.
@@
Given a UFC $\mathcal{C}$, the string-net model built from $\mathcal{C}$ realizes a topological order [15] corresponding to the Drinfeld center $\mathcal{Z}(\mathcal{C})$ [16] which is a well-defined (2+1)D topological quantum field theory (TQFT), or, in mathematical terms, a unitary modular tensor category (UMTC). For example, the vector space dimension assigned to a 2D compact manifold of genus $g$ by $\mathcal{Z}(\mathcal{C})$ matches the ground-state degeneracy of the string-net model built from $\mathcal{C}$ on the same manifold. However, what is often not appreciated is that the agreement between the vector space dimension assigned by $\mathcal{Z}({\mathcal{C}})$ and the degeneracy of states in the string-net model may not be perfect when one considers string-nets on manifolds with boundaries, or the case in which quasiparticles are present, i.e., excitations above the ground state. In particular, for cases other than the ground state on a closed manifold, the string-net model displays both topological and nontopological degeneracies.
@@
By “topological degeneracies,” we mean degeneracies that arise from multiple topological fusion channels in the ${\cal Z}({\cal C})$ topological quantum field theory. As we will discuss below in Se. IV, the degeneracy of the ground state on a surface of nonzero genus is such a topological degeneracy. This type of degeneracy cannot be split by local perturbations as long as the excitation gap does not close. Here, the notion of locality is related to operators acting on a typical scale smaller than the systole (shortest noncontractible loops). Degeneracies that arise from multiple fusion channels of quasiparticle excitations, if present, will also be called topological. These can be split by perturbations only if the perturbing operators connect the positions of the multiple excitations. So for example, if two excitations are very close to each other, a fairly local perturbation operator could in principle split such degeneracy. Topological degeneracies, independent of how easy they are to split, are encoded in the fusion rules of $\mathcal{Z}(\mathcal{C})$.
@@
However, in the case of a string-net model built from a noncommutative UFC there are additional degeneracies associated with quasiparticles that are not topological, and counting them requires knowledge of other quantities that will be discussed below. Indeed, in a given topological sector, one may still have some additional degeneracies that can be lifted by completely local perturbations.
@@
For a string-net model built from a category $\cal C$, certain quasiparticle types of the emergent theory $\mathcal{Z}(\mathcal{C})$ play a special role and are denoted as “fluxons” or plaquette excitations. These elementary excitations are always bosons that can be excited in the string-net model without violating the fusion condition of the model at every trivalent vertex (see Sec. III for more details). In the current paper, we will focus on this type of excitations which only exists for smooth boundaries.
@@
Excitations that are not fluxons and boundaries that are not smooth generally require violations of the fusion condition. This is more complicated since there may be inequivalent ways to define the model once the fusion conditions are violated. Note that in Ref. [17] a particular extension of the string-net model (known as the “extended string-net model”) is considered that contains all topological quasiparticle types but never violates any vertex constraints at the price of some other complications. We leave for future work analysis of that extended model.
@@
It is worth noting that being a fluxon for some particular particle types of a UMTC is dependent on the microscopic string-net model and is not a property of the TQFT. In particular it may happen that the same TQFT may be constructed either as the Drinfeld center $\mathcal{Z}(\mathcal{C})$ of a UFC $\mathcal{C}$ or as the Drinfeld center $\mathcal{Z}(\mathcal{C}^{\prime})$ of another UFC $\mathcal{C}^{\prime}$. When $\mathcal{Z}(\mathcal{C}^{\prime})\simeq\mathcal{Z}(\mathcal{C})$ we say that $\mathcal{C}$ and $\mathcal{C}^{\prime}$ are Morita-equivalent [18]. However, the fluxons of the string-net model built from $\mathcal{C}$ are generally different from the fluxons of the string-net model built from $\mathcal{C}^{\prime}$. This is quite natural since a fluxon is defined as satisfying the fusion rules of the underlying UFC from which the string-net model is built.
@@
In the case in which the fusion rules of the category $\mathcal{C}$ are commutative, i.e., when $a\times b=b\times a$ for all objects $a$ and $b$ in the category, there is a substantial simplification since there are no internal multiplicities. One can then find the topological degeneracy of a 2D orientable manifold with any number of quasiparticles on it using classic formulas found by Moore, Seiberg and Banks [19] and Verlinde [20]. However, in the case in which the fusion rules of the category $\mathcal{C}$ are noncommutative, i.e., $a\times b\neq b\times a$ for some $a$ and $b$ in the category, the situation is more complicated. Here, for some of the fluxons, there is an additional internal degeneracy factor on top of the TQFT/conformal field theory prediction of Moore, Seiberg, Banks, and Verlinde [19, 20]. We will be particularly interested in this case.
@@
Another motivation for the present study is conceptual: it is to clearly understand how much of the physics of the string-net model is captured by the Drinfeld center of the input category and how much is left out. The former may be called topological and the latter non-topological. Here, we will show that beyond the Drinfeld center, one needs quantities called $n_{A,1}$, (see below) extracted from the tube algebra, in order to compute the excited-level degeneracies.
@@
The paper is structured as follows. In Sec. II, we provide a few basic notions on fusion categories. In Sec. III, we give a short introduction to generalized string-net models, with a particular emphasis on the construction and description of the emergent topological phase. Section IV contains our main result, namely the formula for the spectrum degeneracies in the most general case. We make the link with previously obtained results in some particular cases. Special cases such as modular or commutative categories are discussed in Sec. V. We further illustrate our results with examples in Sec. VI. In particular, we study the case of the Morita-equivalent categories Rep($S_{3}$) and Vec($S_{3}$), corresponding to the same Drinfeld center $\mathcal{Z}(S_{3})$ but having different energy-level degeneracies. We also study more involved cases such as the Haagerup, the Tambara-Yamagami, and the Hagge-Hong categories. Eventually, we conclude and discuss perspectives in Sec. VII. In Appendixes, we give details on the tube algebra (see Appendices A and B), on fluxon identities (see Appendix C), on a generalized Hamiltonian (see Appendix D) and on the Hilbert-space dimension (see Appendix E).
@@ II Background information on fusion categories
If there are $N_{\mathcal{C}}$ simple objects in the category, we can think of $N_{ab}^{c}$ as being a set of $N_{\mathcal{C}}$ square matrices $N_{a}$ that are each $N_{\mathcal{C}}$-dimensional, with rows and columns indexed by $b,c$. The quantum dimension $d_{a}$ is defined to be the largest eigenvalue of $N_{a}$, which is guaranteed to be positive. The so-called total quantum dimension is
@@
In a UFC, a fusion vertex, such as the one in Fig. 2, represents a Hilbert space of dimension $N_{ab}^{c}$ with basis vectors labeled with Greek letter $\mu=1\ldots N_{ab}^{c}$. The diagram on the left of Fig. 3 describes a Hilbert space of dimension $\sum_{d}N_{d\bar{b}\bar{a}}N_{e\bar{c}\bar{d}}$ and this space is spanned by the possible values of $d,\mu,\nu$. On the other hand, the diagram on the right describes a Hilbert space of dimension $\sum_{f}N_{f\bar{c}\bar{b}}N_{e\bar{f}\bar{a}}$ and this space is spanned by the possible values of $f,\alpha,\beta$. Due to the associativity condition on fusion, the dimension of the space in the two descriptions are the same, and the unitary $F$-matrix shown in Fig. 3 relates the two bases to each other. There is a consistency condition on the $F$-matrices, known as the pentagon equation [35, 36, 16, 22], that ensures that multiple changes of basis in complicated diagrams will give a consistent result.
@@
[…]
@@
The $S$-matrix is a symmetric matrix of elements $S_{a,b}=S_{b,a}$ that evaluate a loop labeled $a$ linked with a loop labeled $b$ and then normalized with the total quantum dimension, as shown in Fig. 6. It gives the exchange statistics of the anyons $a$ and $b$. In the particular case where one of the anyons is the vacuum, the $S$-matrix element takes a simple value in terms of the quantum dimensions:
@@
A word of caution is in order. Diagrams in Figs. 5 and 6 as well as in Eq. (5) are evaluated, meaning that they give complex numbers corresponding to quantum amplitudes. Most other diagrams that will appear later in this paper, e.g., in Figs. 8 or 9 and in Appendix E, are used differently and as a way to compute the dimension of Hilbert spaces, similarly to what is done in Figs.  2 and  3. The two ways of considering a diagram, i.e., either as a quantum amplitude or as a way of counting possible labelings, should be clearly distinguished. For example, in the second case, loops or bubbles can never be contracted, whereas in the first case, an isolated loop of $a$ is evaluated as the quantum dimension $d_{a}$ [see Eq. (5)].
@@ III Generalized string-net models
Inspired by the original construction proposed by Levin and Wen [10], but without assuming tetrahedral symmetry, the generalized string-net models [12, 11, 14, 13] allow us to generate all topological phases described by the Drinfeld center $\mathcal{Z(C)}$ of a UFC $\mathcal{C}$. To compute the spectral degeneracies of generalized string-net models, we do not have to explain the detailed construction of the model, which can be found in Refs. [14, 13]. We will nonetheless give a rough picture of it here.
The generalized string-net model can be defined on any trivalent graph (see Fig. 1) and its construction only requires an (input) UFC $\mathcal{C}$. The Hilbert space $\mathcal{H}$ is defined by the set of all possible allowed edge and vertex labelings. The trivalent graph is intended to be dual to a triangulation of a genus $g$ two-dimensional orientable manifold.
@@ III.1 The Hamiltonian
The Hamiltonian for a string-net model is written in terms of a set of operators $A_{v}$, defined at the position of each vertex $v$, and operators $B_{p}$ defined at the position of each plaquette $p$. These operators are projectors, meaning that $A_{v}^{2}=A_{v}$ and $B_{p}^{2}=B_{p}$ or equivalently that their eigenvalues are 0 and 1 only. Further, these operators all commute with each other: $[A_{v},A_{v^{\prime}}]=[A_{v},B_{p}]=[B_{p},B_{p^{\prime}}]=0$, for all $v,v^{\prime},p,p^{\prime}$.
@@
with $(J_{v},J_{p})>0$. The $A_{v}$ operators yield unity if the fusion rules of the input category $\mathcal{C}$ are obeyed at vertex $v$, and they yield zero otherwise. In a lattice gauge theory analogy, $A_{v}=1$ corresponds to satisfying a Gauss law at vertex $v$, while $A_{v}=0$ corresponds to having a “charge” at this vertex, i.e., a vertex excitation. The $B_{p}$ operators can be thought of as assuring that no “flux” penetrates the plaquette $p$ (the notion of flux being dual to that of charge). The eigenvalue $B_{p}=1$ corresponds to having no flux in the plaquette $p$ and $B_{p}=0$ corresponds to having a flux, i.e., a plaquette excitation. The detailed form of the $B_{p}$ operator is a bit complicated, and it is given in terms of $F$-symbols of the input UFC. Since we do not need these for our purpose, we refer the interested reader to Refs. [13, 14] for more details (see Refs. [10, 22] for an introduction to the simple case with only commutative fusion rules).

### v4

@@ Abstract
Generalized string-net models have been recently proposed in order to enlarge the set of possible topological quantum phases emerging from the original string-net construction. In the present work, we do not consider vertex excitations and restrict to plaquette excitations, or fluxons, that satisfy important identities. We explain how to compute the energy-level degeneracies of the generalized string-net Hamiltonian associated to an arbitrary unitary fusion category. In contrast to the degeneracy of the ground state, which is purely topological, that of excited energy levels depends not only on the Drinfeld center of the category, but also on internal multiplicities obtained from the tube algebra defined from the category. For a noncommutative category, these internal multiplicities result in extra nontopological degeneracies. Our results are valid for any trivalent graph and any orientable surface. We illustrate our findings with nontrivial examples.
@@ I Introduction
The idea that two-dimensional quantum systems could contain quasiparticles with exchange statistics different from bosons and fermions emerged about fifty years ago [1, 2]. These point-like particles, called anyons, could, under mutual exchange, accumulate “any”, possibly fractional, phase (Abelian anyons), or even change into a new state orthogonal to the initial one (non-Abelian anyons). Yet, only recent experiments provided evidence for the presence of Abelian anyons in the fractional quantum Hall effect [3, 4], and investigated non-Abelian anyons in quantum processors made of superconducting qubits [5, 6] or trapped ions [7]. Anyons are a fundamental characteristic of topologically-ordered phases: gapped phases that appear in strongly interacting or frustrated systems and that cannot be identified by local symmetries. In particular, anyons are connected to one of the defining properties of topological order: a ground-state degeneracy which depends on the surface topology on which the system is defined, and which is robust against small local perturbations (a fact which motivated the idea of using topological order to implement fault tolerant quantum computation  [8, 9]). The excited states of topologically-ordered phases also show nontrivial degeneracies related to the fusion of anyons.
Some of the most studied models for topologically-ordered phases are the string-net models introduced by Levin and Wen [10], and their generalizations [11, 12, 13, 14]. These models, while not being able to produce all possible topological orders, are believed to produce exactly those bosonic topological orders that have gappable edges (and are therefore achiral). The string-net models are exactly solvable. While the energy levels are easy to obtain, the corresponding spectrum degeneracies are nontrivial and are the main focus of the present paper.
@@
Given a UFC $\mathcal{C}$, the string-net model built from $\mathcal{C}$ realizes a topological order [15] corresponding to the Drinfeld center $\mathcal{Z}(\mathcal{C})$ [16] which is a well-defined (2+1)D topological quantum field theory (TQFT), or, in mathematical terms, a unitary modular tensor category (UMTC). For example, the vector space dimension assigned to a 2D compact manifold of genus $g$ by $\mathcal{Z}(\mathcal{C})$ matches the ground-state degeneracy of the string-net model built from $\mathcal{C}$ on the same manifold. However, what is often not appreciated is that the agreement between the vector space dimension assigned by $\mathcal{Z}({\mathcal{C}})$ and the degeneracy of states in the string-net model may not be perfect when one considers string-nets on manifolds with boundary, or the case where quasiparticles are present, i.e., excitations above the ground state. In particular, for cases other than the ground state on a closed manifold, the string-net model displays both topological and nontopological degeneracies.
@@
By “topological degeneracies,” we mean degeneracies that arise from multiple topological fusion channels in the ${\cal Z}({\cal C})$ topological quantum field theory. As we will discuss below in section IV, the degeneracy of the ground state on a surface of nonzero genus is such a topological degeneracy. This type of degeneracy cannot be split by local perturbations so long as the excitation gap does not close. Here, the notion of locality is related to operators acting on a typical scale smaller than the systole (shortest noncontractible loops). Degeneracies that arise from multiple fusion channels of quasiparticle excitations, if present, will also be called topological. These can be split by perturbations only if the perturbing operators connect the positions of the multiple excitations. So for example, if two excitations are very close to each other, a fairly local perturbation operator could in principle split such degeneracy. Topological degeneracies, independent of how easy they are to split, are encoded in the fusion rules of $\mathcal{Z}(\mathcal{C})$.
@@
However, in the case of a string-net model built from a noncommutative UFC there are additional degeneracies associated with quasiparticles that are not topological, and counting them requires the knowledge of other quantities that will be discussed below. Indeed, in a given topological sector, one may still have some additional degeneracies which can be lifted by completely local perturbations.
@@
For a string-net model built from a category $\cal C$, certain quasiparticle types of the emergent theory $\mathcal{Z}(\mathcal{C})$ play a special role and are denoted as “fluxons” or plaquette excitations. These elementary excitations are always bosons that can be excited in the string-net model without violating the fusion condition of the model at every trivalent vertex (see Sec. III for more details). In the current paper, we will focus on this type of excitations.
@@
Excitations that are not fluxons and boundaries that are not smooth generally require violations of the fusion condition. This is more complicated since there may be inequivalent ways to define the model once the fusion conditions are violated. Note that in Ref. [17] a particular extension of the string-net model (known as “extended string-net model”) is considered that contains all topological quasiparticle types but never violates any vertex constraints at the price of some other complications. We leave for future work analysis of that extended model.
@@
It is worth noting that being a fluxon for some particular particle types of a UMTC is dependent on the microscopic string-net model and is not a property of the TQFT. In particular it may happen that the same TQFT may be constructed either as the Drinfeld center $\mathcal{Z}(\mathcal{C})$ of a UFC $\mathcal{C}$ or as the Drinfeld center $\mathcal{Z}(\mathcal{C}^{\prime})$ of another UFC $\mathcal{C}^{\prime}$. When $\mathcal{Z}(\mathcal{C}^{\prime})\simeq\mathcal{Z}(\mathcal{C})$ we say that $\mathcal{C}$ and $\mathcal{C}^{\prime}$ are Morita equivalent [18]. However, the fluxons of the string-net model built from $\mathcal{C}$ are generally different from the fluxons of the string-net model built from $\mathcal{C}^{\prime}$. This is quite natural since a fluxon is defined as satisfying the fusion rules of the underlying UFC that the string-net model is built from.
@@
In the case where the fusion rules of the category $\mathcal{C}$ are commutative, i.e., when $a\times b=b\times a$ for all objects $a$ and $b$ in the category, there is a substantial simplification since there are no internal multiplicities. One can then find the topological degeneracy of a 2D orientable manifold with any number of quasiparticles on it using classic formulas found by Moore, Seiberg and Banks [19] and Verlinde [20]. However, in the case where the fusion rules of the category $\mathcal{C}$ are noncommutative, i.e., $a\times b\neq b\times a$ for some $a$ and $b$ in the category, the situation is more complicated. Here, for some of the fluxons, there is an additional internal degeneracy factor on top of the TQFT/conformal field theory prediction of Moore, Seiberg, Banks, and Verlinde [19, 20]. We will be particularly interested in this case.
@@
Another motivation for the present study is conceptual: it is to clearly understand how much of the physics of the string-net model is captured by the Drinfeld center of the input category and how much is left out. The former may be called topological and the latter nontopological. Here, we will show that beyond the Drinfeld center, one needs quantities called $n_{A,1}$, (see below) extracted from the tube algebra, in order to compute the excited-level degeneracies.
@@
The paper is structured as follows. In Sec. II, we provide a few basic notions on fusion categories. In Sec. III, we give a short introduction to generalized string-net models, with a particular emphasis on the construction and description of the emergent topological phase. Section IV contains our main result, namely the formula for the spectrum degeneracies in the most general case. We make the link with previously obtained results in some particular cases. Special cases such as modular or commutative categories are discussed in Sec. V. We further illustrate our results with examples in Sec. VI. In particular, we study the case of the Morita-equivalent categories Rep($S_{3}$) and Vec($S_{3}$), corresponding to the same Drinfeld center $\mathcal{Z}(S_{3})$ but having different energy-level degeneracies. We also study more involved cases such as the Haagerup, the Tambara-Yamagami and the Haage-Hong categories. Eventually, we conclude and discuss perspectives in Sec. VII. In appendices, we give details on the tube algebra (see Appendices A and B), on fluxon identities (see Appendix C), on a generalized Hamiltonian (see Appendix D) and on the Hilbert-space dimension (see Appendix E).
@@ II Background information on fusion categories
If there are $N_{\mathcal{C}}$ simple objects in the category, we can think $N_{ab}^{c}$ as being a set of $N_{\mathcal{C}}$ square matrices $N_{a}$ which are each $N_{\mathcal{C}}$-dimensional, with rows and columns indexed by $b,c$. The quantum dimension $d_{a}$ is defined to be the largest eigenvalue of $N_{a}$, which is guaranteed to be positive. The so-called total quantum dimension is
@@
In a UFC, a fusion vertex, such as the one in Fig. 2, represents a Hilbert space of dimension $N_{ab}^{c}$ with basis vectors labeled with greek letter $\mu=1\ldots N_{ab}^{c}$. The diagram on the left of Fig. 3 describes a Hilbert space of dimension $\sum_{d}N_{d\bar{b}\bar{a}}N_{e\bar{c}\bar{d}}$ and this space is spanned by the possible values of $d,\mu,\nu$. On the other hand, the diagram on the right describes a Hilbert space of dimension $\sum_{f}N_{f\bar{c}\bar{b}}N_{e\bar{f}\bar{a}}$ and this space is spanned by the possible values of $f,\alpha,\beta$. Due to the associativity condition on fusion, the dimension of the space in the two descriptions are the same, and the unitary $F$-matrix shown in Fig. 3 relates the two bases to each other. There is a consistency condition on the $F$-matrices, known as the pentagon equation [35, 36, 16, 22], that assures that multiple changes of basis in complicated diagrams will give a consistent result.
@@
When fusion rules are commutative, several types of UFC can further be distinguished as depicted in Fig. 4. First, one may ask whether a commutative UFC is braided or not. A unitary braided fusion category (UBFC) is a UFC obeying the so-called hexagon equations [35, 36, 16, 22]. These equations stems from consistency condition required when extending the planar diagram rules to diagrams with over- and under-crossings. Thus (2+1)-dimensional diagrams, corresponding essentially to space-time diagrams for world-lines of anyons, can be given values. The rules for handling over- and under-crossings (known as $R$-matrix) are described by these hexagon equations. It may be the case that for a given UFC, there is no consistent solution to these equations, and the UFC is not braided. For instance, the Tambara-Yamagami category TY3 [37], which we will discuss further below, is commutative but non braided.
@@
The $S$-matrix is a symmetric matrix of elements $S_{a,b}=S_{b,a}$ which evaluate a loop labeled $a$ linked with a loop labeled $b$ and then normalized with the total quantum dimension, as shown in Fig. 6. It gives the exchange statistics of the anyons $a$ and $b$. In the particular case where one of the anyons is the vacuum, the $S$-matrix element takes a simple value in terms of the quantum dimensions:
@@
A word of caution is in order. Diagrams in Figs. 5 and 6 as well as in Eq. (5) are evaluated meaning that they give complex numbers corresponding to quantum amplitudes. Most other diagrams that will appear later in this paper, e.g., in Figs. 8 or 9 and in Appendix E, are used differently and as a way to compute the dimension of Hilbert spaces, similarly to what is done in Figs.  2 and  3. The two ways of considering a diagram, i.e., either as a quantum amplitude or as a way of counting possible labelings, should be clearly distinguished. For example, in the second case, loops or bubbles can never be contracted, whereas in the first case, an isolated loop of $a$ is evaluated as the quantum dimension $d_{a}$ [see Eq. (5)].
@@ III Generalized string-net models
Inspired by the original construction proposed by Levin and Wen [10], but without assuming tetrahedral symmetry, the generalized string-net models [12, 11, 14, 13] allow us to generate all topological phases described by the Drinfeld center $\mathcal{Z(C)}$ of a UFC $\mathcal{C}$. To compute the spectral degeneracies of generalized string-net models, we do not have to explain the detailed construction of the model which can be found in Refs. [14, 13]. We will nonetheless give a rough picture of it here.
The generalized string-net model can be defined on any trivalent graph (see Fig. 1) and its construction only requires a (input) UFC $\mathcal{C}$. The Hilbert space $\mathcal{H}$ is defined by the set of all possible allowed edge and vertex labelings. The trivalent graph is intended to be dual to a triangulation of a genus $g$ two-dimensional orientable manifold.
@@ III.1 The Hamiltonian
The Hamiltonian for a string-net model is written in terms of a set of operators $A_{v}$, defined at the position of each vertex $v$, and operators $B_{p}$ defined at the position of each plaquette $p$. These operators are projectors, meaning that $A_{v}^{2}=A_{v}$ and $B_{p}^{2}=B_{p}$ or equivalently that their eigenvalues are 0 and 1 only. Further these operators all commute with each other: $[A_{v},A_{v^{\prime}}]=[A_{v},B_{p}]=[B_{p},B_{p^{\prime}}]=0$, for all $v,v^{\prime},p,p^{\prime}$.
@@
with $(J_{v},J_{p})>0$. The $A_{v}$ operators yield unity if the fusion rules of the input category $\mathcal{C}$ are obeyed at vertex $v$, and give zero otherwise. In a lattice gauge theory analogy, $A_{v}=1$ corresponds to satisfying a Gauss law at vertex $v$, while $A_{v}=0$ corresponds to having a “charge” at this vertex, i.e., a vertex excitation. The $B_{p}$ operators can be thought of as assuring that no “flux” penetrates the plaquette $p$ (the notion of flux being dual to that of charge). The eigenvalue $B_{p}=1$ corresponds to having no flux in the plaquette $p$ and $B_{p}=0$ to having a flux, i.e., a plaquette excitation. The detailed form of the $B_{p}$ operator is a bit complicated, and is given in terms of $F$-symbols of the input UFC. Since we do not need these for our purpose, we refer the interested reader to Refs. [13, 14] for more details (see Refs. [10, 22] for an introduction to the simple case with only commutative fusion rules).